a. From the information given here determine the
95% confidence interval estimate of the population
mean.
x = 100  = 20 n = 25
b. Repeat part (a) with x = 200.
c. Repeat part (a) with x = 500.
d. Describe what happens to the width of the confidence
interval estimate when the sample mean
increases.
1
Expert's answer
2020-12-15T02:15:19-0500

95 % CI="\bar x±Z_\frac{\alpha}{2}\frac{\sigma}{\sqrt n}"

a) "\bar x=100"

="100±1.96* \frac {20}{\sqrt{25}}"

=(92.16, 107.84)

b)"\bar x=200"

"=200±1.96* \frac {20}{\sqrt{25}}"

=(192.16, 207.84)

c) "\bar x=500"

="500±1.96* \frac {20}{\sqrt{25}}"

=(492.16, 507.84)

d) As the sample mean increases the confidence interval width remains constant. This is because the standard error is constant for all sample means.